# DRESS and the WL Hierarchy: Climbing One Deletion at a Time

### v2

@@ 4.2 Analysis
    $\Delta^{0}$ (Original-DRESS) distinguishes $\text{CFI}(K_{3})$, which requires 2-WL. This is a notable result: Original-DRESS—with no node deletion—already exceeds 1-WL on CFI instances, consistent with the prism vs. $K_{3,3}$ result in [4]. It fails on all larger CFI pairs.
@@
    $\Delta^{3}$ reaches $\text{CFI}(K_{6})$ (5-WL) but fails on $\text{CFI}(K_{7})$ (6-WL), confirming that the staircase has a sharp boundary at $(\ell+2)$-WL. The $\text{CFI}(K_{7})$ test required processing all $\binom{224}{3}=1{,}848{,}224$ subgraphs ($\approx$19.3 billion edge values).
@@
Deletion depth $\ell$ Max WL matched Effective WL
0 2-WL $\ell+2$
1 3-WL $\ell+2$
2 4-WL $\ell+2$
3 5-WL $\ell+2$
The pattern is strikingly regular: $\Delta^{\ell}$-DRESS empirically matches $(\ell+2)$-WL on the CFI family. Each deletion level adds one WL level of expressiveness, with the base DRESS contributing 2-WL on its own. The computational cost grows as $\binom{n}{\ell}$ (polynomial in $n$ for fixed $\ell$), while the equivalent $(\ell+2)$-WL cost grows as $n^{\ell+3}$.
@@ 5.1 Relationship to Existing Work
Methods such as ESAN [1] and GNN-AK+ [8] also use node-deleted or node-marked subgraphs to boost expressiveness. However, these are supervised methods that learn aggregation functions from data. $\Delta^{\ell}$-DRESS is entirely unsupervised: the aggregation is the deterministic DRESS fixed point, and the histogram comparison is parameter-free. This makes $\Delta^{\ell}$-DRESS a canonical baseline for the expressiveness of subgraph-based approaches.
@@
The Kelly–Ulam conjecture [5, 6] states that graphs with $n\geq 3$ are determined by their deck of node-deleted subgraphs. Conceptually, $\Delta^{1}$-DRESS computes a continuous relaxation of this deck: each “card” is the DRESS fingerprint of a node-deleted subgraph. In practice, all edge values are accumulated into a fixed-size histogram for memory efficiency, though individual cards are not recoverable.
@@
For fixed $\ell$, $\Delta^{\ell}$-DRESS runs in polynomial time. The empirical pattern $\Delta^{\ell}\approx(\ell+2)$-WL suggests that the minimum deletion depth needed to distinguish $\text{CFI}(K_{n})$ grows as n2n-2, which is linear in $n$. While each additional deletion level is polynomial, the number of required levels grows with the base graph size.
@@ 5.2 Open Questions
    Does the $\Delta^{\ell}\approx(\ell+2)$-WL pattern continue? We have confirmed that $\Delta^{3}$ fails on $\text{CFI}(K_{7})$ (6-WL). Testing $\Delta^{4}$ on $\text{CFI}(K_{7})$ would confirm whether the staircase extends to $\ell=4$.
@@
    Is the pattern specific to CFI, or does it hold more broadly? Other hard graph families may behave differently.
@@
    Are there graph families where $\Delta^{\ell}$-DRESS fails for all constant $\ell$? CFI graphs over graphs of unbounded treewidth would be natural candidates.
@@
    Does combining $\Delta^{\ell}$ with other DRESS variants yield further gains? The DRESS family includes Motif-DRESS and Generalized-DRESS [4]. Whether a stronger base variant $\mathcal{F}$ shifts the staircase upward or merely overlaps with the deletion mechanism is an open question.
@@ 6 Conclusion
We introduced $\Delta^{\ell}$-DRESS, which applies $\ell$ levels of iterated node deletion to the DRESS framework. Original-DRESS ($\Delta^{0}$) already distinguishes $\text{CFI}(K_{3})$, and each additional deletion level adds one WL level: $\Delta^{1}$ matches 3-WL, $\Delta^{2}$ matches 4-WL, and $\Delta^{3}$ matches 5-WL but fails on 6-WL, establishing the empirical pattern $\Delta^{\ell}\approx(\ell+2)$-WL on the CFI family with a sharp upper boundary. These results establish $\Delta^{\ell}$-DRESS as a systematic, scalable method for climbing the WL hierarchy on the canonical CFI benchmarks. An open-source implementation is available at https://github.com/velicast/dress-graph.

### v3

@@ 5.3 Analysis
    $\Delta^{0}$ (Original-DRESS) distinguishes $\text{CFI}(K_{3})$, which requires 2-WL. This is a notable result: Original-DRESS–with no vertex deletion–already exceeds 1-WL on CFI instances, consistent with the prism vs. $K_{3,3}$ result in [4]. It fails on all larger CFI pairs.
@@
    $\Delta^{3}$ reaches $\text{CFI}(K_{6})$ (5-WL) but fails on $\text{CFI}(K_{7})$ (6-WL), confirming that the staircase has a sharp boundary at $(k+2)$-WL. The $\text{CFI}(K_{7})$ test required processing all $\binom{224}{3}=1{,}848{,}224$ subgraphs ($\approx$19.3 billion edge values).
@@
Deletion depth $k$ Max WL matched Effective WL
0 2-WL $k+2$
1 3-WL $k+2$
2 4-WL $k+2$
3 5-WL $k+2$
The pattern is strikingly regular: $\Delta^{k}$-DRESS empirically matches $(k+2)$-WL on the CFI family, consistent with the lower bound proved in Theorem 8. Each level adds one WL level of expressiveness, with the base DRESS contributing 2-WL on its own. The computational cost grows as $\binom{n}{k}$ (polynomial in $n$ for fixed $k$), while the equivalent $(k+2)$-WL cost grows as $n^{k+3}$.
@@ 6.1 Relationship to Existing Work
Methods such as ESAN [1] and GNN-AK+ [10] also use vertex-deleted or vertex-marked subgraphs to boost expressiveness. However, these are supervised methods that learn aggregation functions from data. $\Delta^{k}$-DRESS is entirely unsupervised: the aggregation is the deterministic DRESS fixed point, and the histogram comparison is parameter-free. $\Delta^{k}$-DRESS corresponds directly to the vertex-deletion approach used in subgraph GNNs.
@@
The Kelly–Ulam conjecture [7, 8] states that graphs with $n\geq 3$ are determined by their deck of vertex-deleted subgraphs. Conceptually, $\Delta^{1}$-DRESS computes a continuous relaxation of this deck: each “card” is the DRESS fingerprint of a vertex-deleted subgraph. The expressiveness proof (Theorem 8) relies on the Reconstruction Conjecture to bridge vertex deletion and the WL individualization hierarchy. The empirical success of $\Delta^{k}$-DRESS on all tested CFI instances is consistent with RC, and provides further evidence for the conjecture on these graph families.
@@
For fixed $k$, $\Delta^{k}$-DRESS runs in polynomial time. Theorem 8 guarantees (under RC) that $\Delta^{k}$-DRESS is at least as powerful as $(k+2)$-WL, and the CFI experiments confirm this bound is tight: the minimum depth needed to distinguish $\text{CFI}(K_{n})$ grows as n2n-2, which is linear in $n$. While each additional level is polynomial, the number of required levels grows with the base graph size.
@@ 6.2 Open Questions
    Does the staircase pattern continue beyond $k=3$? Theorem 8 establishes $\geq(k+2)$-WL (under RC). We have confirmed empirically that $k=3$ fails on $\text{CFI}(K_{7})$ (6-WL), so the bound is tight for $k\leq 3$. Testing $k=4$ on $\text{CFI}(K_{7})$ would confirm whether the staircase extends.
@@
    Is the tightness specific to CFI, or does it hold more broadly? The lower bound $(k+2)$-WL is general, but the upper bound may depend on the graph family.
@@
    Are there graph families where $\Delta^{k}$-DRESS fails for all constant $k$? CFI graphs over graphs of unbounded treewidth would be natural candidates.
@@
    Does combining iterated deletion with other DRESS variants yield further gains? The DRESS family includes Motif-DRESS and Generalized-DRESS [4]. Whether a stronger base variant $\mathcal{F}$ shifts the staircase upward or merely overlaps with the deletion mechanism is an open question.
  5. 5.
    Can the RC assumption be removed? Theorem 8 relies on the Reconstruction Conjecture. An interesting direction is whether the expressiveness guarantee can be proved unconditionally, perhaps via alternative structural arguments.
@@ 7 Conclusion
We introduced $\Delta^{k}$-DRESS, which applies $k$ levels of iterated vertex deletion to the DRESS framework. We proved that $\Delta^{k}$-DRESS is at least as powerful as $(k+2)$-WL under the Kelly–Ulam Reconstruction Conjecture (Theorem 8): the reconstruction deck bridges vertex deletion and the WL individualization hierarchy, allowing the inductive argument to lift one WL level per deletion depth. $\Delta^{k}$-DRESS was verified empirically on the canonical CFI benchmark family: at $k=0$, Original-DRESS already distinguishes $\text{CFI}(K_{3})$, and each additional level adds one WL level. $k=3$ reaches 5-WL but fails on 6-WL, confirming the bound is tight on CFI instances. These results establish iterated DRESS as a systematic, scalable method for climbing the WL hierarchy with provable expressiveness guarantees (modulo RC). An open-source implementation is available at https://github.com/velicast/dress-graph.